THE CORONENE VIBRONIC STATES ABOVE THE FIRST IONIZATION POTENTIAL INVESTIGATED THROUGH TPEPICO EXPERIMENTS

\underline{PH. BRECHIGNAC}, C. FALVO, P. PARNEIX , T. PINO, O. PIRALI, Institut des Sciences Moléculaires d’Orsay, CNRS UMR8214, Univ Paris-Sud, Bât 210, F91405 Orsay Cedex, France; G. GARCIA, L. NAHON, Synchrotron SOLEIL, L’ Orme des Merisiers, St Aubin, B.P. 48, 91192 Gif sur Yvette, France; C. JOBLIN, D. KOKKIN, A. BONAMMY, IRAP, Université de Toulouse [UPS], CNRS, Toulouse, France; G. MULAS, INAF-Osservatorio Astronomico di Cagliari-Astrochemistry Group, Strada 54, Località Poggio dei Pini, I-09012 Capoterra (CA), Italy.

Threshold Photoelectron spectra (TPES), as well as Total Ion Yeld (TIY) spectra of jet-cooled Coronene (C_{12}H_{24}) have been obtained using the electron/ion coincidence imaging spectrometer DELICIOUS II^{a} available at the DESIRS beamline of the French Synchrotron facility SOLEIL. The obtained data can be interpreted in the light of new theoretical DFT and TDDFT based calculations. They will be discussed in comparison to available photoelectron (PES) and optical absorption data. New autoionizing neutral states have also been observed and their relaxation to final cationic states characterized.

aG.A. Garcia, H. Soldi-Lose, and L. Nahon, Rev. Sci. Instrum. 80, 023102 (2009).